A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 183-cm and a standard deviation of 0.6-cm. For shipment, 20 steel rods are bundled together. Find the probability that the average length of a randomly selected bundle of steel rods is less than 183.1-cm.

P( ¯ x < 183.1-cm) =______

Enter your answer as a number accurate to 4 decimal places.
Answer #1

let X be length of steel rod x ny Ne u =183.8=0.6) here n=20 we know that if x D N1418) then Xs N(4= 4, 6 = 6/W) ex = 183 byP (X< 183)) = P(Z < 183.1-प- 65 पत =p (२ 183-83 1342 =P (2 <0.7452) - 0.7719 .. P(X< 183. Cm): 0.771g